OpenStudy (anonymous):

True or false: When a bullet is fired horizontally, it takes the same amount of time to reach the ground as a bullet dropped from rest from the same height. Why?

OpenStudy (bahrom7893):

True

OpenStudy (bahrom7893):

Because the initial vertical component of a bullet's velocity is 0 in both cases

OpenStudy (anonymous):

Officially:"The x and y components of the motion of ANYTHING is independent" What does this MEAN? Take any particle or body or whatever and apply a force onto it with x and y components. The force and resulting motion in the x direction is exactly the same if you just applied the x component of that force in the x direction. It's like 2 different worlds are in motion at once.

OpenStudy (anonymous):

Of course in 3D, that would be 3 worlds.

OpenStudy (bahrom7893):

Basically if you drop the bullet, then Vi = 0 Viy = ViSin(theta) = 0*Sin(theta) = 0 If you shoot it horizontally, then theta=0, so Viy = ViSin(theta) = Vi * Sin0 = Vi* 0 = 0

OpenStudy (anonymous):

so in the Y direction of motion there is only gravity acting on the bullet, in the first case, the velocity is inthe X direction so its only important for the Range but not the height and the gravity alone acts on the bullet on hte Y axis, and in hte free fall gravity os the only thing acting on it regardless if there is motion in the x or not

OpenStudy (anonymous):

Or: if you projected a shadow from the x-y plane onto just the x plane, you wouldn't be able to tell (with forces and momentum) what was going on in the y-direction (although in practice you could make a guess, as the 'energy' of the shadows aren't conserved, but oh well)

OpenStudy (anonymous):

so True of course!!!

OpenStudy (anonymous):

Hmm... Interesting. Thanks for the answers!